Problem 32.12 A capacitor has a peak current of 310 μA when the peak voltage at 340 kHz is 2.5 V Part A What is the capacitance? Express your answer to two significant figures and include the appropriate units. c= Value Units Submit Give Up Part B If the peak voltage is held constant, what is the peak current at 680 kHz ? Express your answer to two significant figures and include the appropriate units. le=1 Value Units
Answer #1

a) Given

     Imax = 310*10^-6 A

     Vmax = 2.5 V

But we have Xc = Vmax/Imax = 2.5/(310*10^-6) = 8064.52 ohm

Xc = 1/(2pi*f*C) =======> C = 1/(2pi*f*Xc) = 1/(2pi*340000*8064.52)

C = 58*10^-12 F = 58 pF

(b) Vmax = 2.5 V

      f = 680000 Hz

      C = 58*10^-12 F

      Xc = 1/(2pi*f*C) = 4035.4 ohm

       Imax = Vmax/Xc = 2.5/4035.4 = 6.195*10^-4 A = 619.5 uA = 58*10^-12 F

       Xc = 1/(2pi*f*C) = 4035.4 ohm

Now Imax = Vmax/Xc = 2.5/4035.4 = 619.5 uA